OpenStudy (anonymous):

What is the arc length of the subtending arc for an angle of 72 degrees on a circle of radius 4?

OpenStudy (johnweldon1993):

Hint \(\large s = r\theta \) But you need \( \theta\) in radians to measure

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

You are looking for s.

OpenStudy (mathstudent55):

The arc length of an entire circle is its circumference. Do you know how to find the circumference of a circle of radius 4?

OpenStudy (mathstudent55):

Find the circumference of the circle. That is the arc length of an arc tended by a 360-deg central angle. Then use a proportion to find what the arc length is for a central angle of 72 deg. You can also use this formula: \(s = \dfrac{n}{360^o}\cdot 2 \pi r\) or simply \(s = \dfrac{n \pi r}{180^o}\) where n is the measure of the central angle.